# RLHF Workflow: From Reward Modeling to Online RLHF

### v2

@@ 2 Reward Modeling as Human Feedback Approximation
We present the details of preference model construction in this section, where we study both the reward modeling as MLE of the BT model and the general preference model. The training script and full recipe are provided in https://github.com/RLHFlow/RLHF-Reward-Modeling.
HH-RLHF 115092 160.4 82.2 73.6 LLM Human 115396
SHP 31003 186.2 173.6 88.8 Human Human 93301
HelpSteer 8592 530 116.4 89.3 LLM Human 37131
PKU-SafeRLHF-30K 6975 21.5 70.4 74.6 LLM Human 26874
UltraFeedback 63591 161.5 279.5 211.1 LLM GPT-4 340025
UltraInteract 76086 507.4 396.6 416.7 LLM GPT-4 161927
CodeUltraFeedback 9938 172.8 427.6 400.6 LLM GPT-3.5 50156
Argilla-Math 2352 36.5 276.5 265.3 LLM GPT-4 2418
OpenOrca 6791 153.3 165.4 260.5 LLM GPT-4 6926
Capybara 14740 634.5 348.4 401.9 LLM GPT-4 14811
2.1 Preference Datasets
Following the Pair-RM (Jiang et al.,, 2023) and LLaMA-2 (Touvron et al.,, 2023), we use a mixture of open-source datasets as the training set. Here is a brief introduction to the datasets:
  • HH-RLHF (Bai et al., 2022a, ) is a pairwise preference dataset where each sample is accompanied by a conversation history and two alternative responses written by an early Claude model with 52B parameters. The preferences of the responses are annotated by humans.
  • SHP (Ethayarajh et al.,, 2022) is sourced from Reddit and includes examples from 18 subreddits, such as askacademia, askbaking, askengineers, and changemyview. Each example is a Reddit post with a question/instruction and a pair of top-level comments. One comment is preferred by more Reddit users than the other. All preferences and responses are provided by humans. Only samples with a score ratio $>$ 2 are used, and at most 5 pairs are taken for each prompt.
  • HelpSteer (Wang et al.,, 2023). This open-source dataset (Wang et al.,, 2023) contains prompts, responses, and five human-annotated attributes (helpfulness, correctness, coherence, complexity, and verbosity) ranging from 0 to 4. The prompts are generated using a mixture of template-generated and human-generated methods, while responses are generated by an in-house LLM. The authors generate up to 4 responses per prompt, and we can construct pairwise comparisons based on them.
  • PKU-SafeRLHF (Ji et al.,, 2024). This dataset (Ji et al.,, 2024) consists of 30k+ expert comparison data. Each sample includes two responses to a question and two preference signals for helpfulness and safety, respectively. The responses are generated by open-source chatbots, and the preference signals are merged through the results of 14 harm category multi-class classficiation.
  • UltraFeedback (Cui et al.,, 2023) consists of 64k prompts from diverse resources (including UltraChat, ShareGPT, Evol-Instruct, TruthfulQA, FalseQA, and FLAN) and the authors generate 4 responses per prompt using 4 different LLMs sampled from a diverse set of state-of-the-art open-source LLMs. The preference is from GPT-4 based on a fine-grained annotation instruction, which contains 4 different aspects, namely instruction-following, truthfulness, honesty and helpfulness. The dataset collection strategy of UltraFeedback has also influenced many subsequent works.
  • UltraInteract (Yuan et al., 2024a, ) is a preference dataset designed for complex reasoning tasks. The authors collect a preference tree for each instruction, with the instruction being the root and each action a node. A trajectory is a root-to-leaf path consisting of a sequence of actions. Paired correct and incorrect nodes or trajectories are used for preference learning.
  • Distilabel-Capybara [^2] https://huggingface.co/datasets/argilla/distilabel-capybara-dpo-7k-binarized is a preference dataset of multi-turn dialogues whose prompts are taken from Daniele and Suphavadeeprasit, (2023), where the responses are generated by open-source LLMs and preferences are generated by GPT-4.
  • Distilabel-Orca [^3] https://huggingface.co/datasets/argilla/distilabel-intel-orca-dpo-pairs is collected similarly with Capybara but with the prompts from Lian et al., 2023a .
The training of LLMs is highly data-dependent. To ensure high-quality training data, we conduct a filtering process on the open-source datasets we use. This process removes low-quality and meaningless samples. Additionally, conversations with empty rounds or incorrect labels (implied by the other features of the dataset) are eliminated. Furthermore, in datasets where absolute scores are available, pairwise comparisons with small margins are excluded as these preference signals tend to be noisy (Bansal et al.,, 2023). This process roughly deletes 10% of the data. We summarize the statistics of the open-source datasets that are used for the training in Table 1 and prepare them, as well as our data filtering script, on the huggingface, which is available at https://huggingface.co/collections/RLHFlow/standard-format-preference-dataset-662eec0252e194d5d40c252a.
@@ 2.2 Bradley-Terry Reward Model and Preference Model
Bradley-Terry model construction. We follow the previous works (Ouyang et al.,, 2022; Bai et al., 2022a, ) to initialize the reward model using the SFT model [^4] For preference/reward modeling, we use meta-llama/Meta-Llama-3-8B-Instruct since only this checkpoint is available in the early stage of this project. . We replace the last layer with a linear head to predict a scalar score suitable for preference learning. The reward model is trained using the negative log-likelihood loss function, enabling maximum likelihood estimation (MLE). This loss function is defined as:
@@
where $a^{w}$ is the preferred response over $a^{l}$. We train the LLaMA-3-8B-based reward model for one epoch with a global batch size of 512. The learning rate is set to $\mathrm{lr}=2\times 10^{-6}$, and a cosine learning rate schedule with a warm-up ratio of 0.03 is employed. For Gemma-2B-it, a larger learning rate of $10^{-5}$ is used.
Preference model construction. A (pairwise) preference model takes a prompt $x$ and two responses $a^{1},a^{2}$ as the input and predicts the probability of ^$\hat{\mathbb{P}}(a^{1}\succ a^{2}|x,a^{1},a^{2})$ (Jiang et al.,, 2023). We follow Zhao et al., (2023); Liu et al., 2023a to leverage the LLM’s capability as a next-token predictor for preference modeling. Specifically, for a given preference pair $(x,a^{1},a^{2},A)$, where $A$ indicates that the first response is preferred, the pair is formatted as an instruction-following task:
@@
If the second response is preferred, we replace the label A with B. Then, we simply treat the preference modeling as an instruction-following task to fine-tune the model on these instruction-label pairs. To mitigate position bias (the preference model may prefer the response that is given in the position of RESPONSE A), the order of the responses is randomized during data formatting. During inference, we simply use the probability of decoding A as the ^$\hat{\mathbb{P}}(a^{1}\succ a^{2}|x,a^{1},a^{2})$.
We train the LLaMA-3-8B-based preference model for one epoch. The samples are packed into blocks with length 3072 and a global batch size of 128 is used. The learning rate is set to $\mathrm{lr}=5\times 10^{-6}$, and a cosine learning rate schedule with a warm-up ratio of 0.03 is employed. For Gemma-2B-it, a larger learning rate of $10^{-5}$ is used. We mention in passing that it is possible to include detailed rubrics in the data format to further improve the preference dataset, which we leave for future work (Qin et al.,, 2023).
@@ 2.3 Evaluation Result
We consider two versions of the training set:
  • Mix1: HH-RLHF + SHP + UltraFeedback + Summarization (Stiennon et al.,, 2020).
  • Mix2: all the datasets in Table 1.
The Mix1 dataset is similar to the construction of UltraFeedback (Cui et al.,, 2023) with an additional summarization dataset. In comparison, the Mix2 consists of more reasoning preference pairs (math and code) and safety data. We also consider three different approaches to model the preference signals, including prompting in the LLM-as-a-judge manner (Zheng et al.,, 2023), reward modeling as the MLE of the BT reward model, and the preference model.
We evaluate the models using the RewardBench (Lambert et al.,, 2024), a benchmark designed to assess reward model capabilities in four categories: Chat, Chat-Hard, Safety, and Reasoning. The main evaluation results are in Table 2. It is evident that without explicit training, the prompting approach is inferior to both the BT model and the preference model across all metrics. Meanwhile, the preference model outperforms the BT model in reasoning tasks related to coding and math. We also notice that with more data, especially data specified in coding, math, and safety, the reward model trained by mix2 achieves higher accuracy in safety and reasoning compared with early versions of attempts. In particular, we use the Ultra-RM-13B as a reference model in Table 2, where we can observe that the extra data related to safety and reasoning (as well as a stronger base model) largely contribute to the superior performance of our reward model.
Length bias in reward modeling. It is known that the LLMs aligned by RLHF usually give longer responses (Xiong et al.,, 2023; Yuan et al., 2024b, ; Yuan et al., 2024b, ), where the length bias also exists in the reward models, likely influenced by the preference data used. To better understand this bias, we randomly sample 2K prompts from the prompt set and use the SFT model to generate 8 responses per prompt (see the details in Section 3). Then, we compute the lengths and rewards of the responses and plot the heatmaps of the Pearson correlation coefficient between them in Figure 5. Clearly, both of the two reward models are biased toward the longer responses to some degree. In comparison, UltraRM-13B demonstrates a stronger bias, as we observe that the mean Pearson correlation coefficient of the UltraRM-13B (left figure) is 0.19, while it is 0.06 for our BT reward trained on mix2 (right figure). This may partially result from the use of additional Capybara, OpenOrca, and UltraInteract, whose preferred responses are shorter than the rejected responses. We will return to the ablation study of the impacts of the reward models in Section 4.
Figure 5: The heatmap of the Pearson correlation coefficients between reward and response length. For each prompt, we use the SFT model to generate 16 responses and compute the coefficient. We also group the prompts by the mean responses (y-axis). The left figure is the UltraRM-13B and the right one is our BT reward based on mix2.
Table 2: Comparison of the test accuracy between the Bradley-Terry (BT) reward model and the preference model. We evaluate the model using the Reward-Bench (Lambert et al.,, 2024).
@@
We mention in passing that in the literature, it is also common to model the different types of signals separately, and strategically merge them in the subsequent policy optimization stage (Touvron et al.,, 2023; Wang et al.,, 2024), which we leave for future work.
@@ 3 Iterative Policy Optimization
We develop the main algorithms for the online iterative RLHF in this section, with both theoretical insights and implementation details. In particular, the algorithms will be designed in a direct preference learning style for stable and efficient training.
@@ 3.1 Supervised Fine-Tuning
The base model used in this project is LLaMA-3-8B. To ensure the reproducibility and openness of the project, we perform SFT by ourselves to obtain the initial policy $\pi_{0}$. We collect a set of high-quality instruction datasets for SFT, such as ShareGPT, SlimOrca (Lian et al., 2023b, ), MathInstruct (Yue et al.,, 2023), and Evol-Instruct (see the Appendix for a full list). The training is carried out for one epoch with a learning rate of $2\times 10^{-5}$. A cosine scheduler is employed, and the global batch size is set to 32 with a warm-up ratio of 0.03. To accelerate training, we follow Diao et al., (2023); Tunstall et al., (2023) to pack the samples and use a block size of 8192. We use the SFT-Trainer from TRL and the detailed recipe is available at https://github.com/RLHFlow/Online-RLHF.
@@ 3.2 Iterative Direct Preference Learning: Theoretical Insights and Algorithmic Principles
Hybrid batch learning. We formulate a slightly more general framework to combine an initial offline dataset with online data collected during training, hence the name hybrid learning, similar to the recipe of Claude (Bai et al., 2022a, ) and LLaMA-2 (Touvron et al.,, 2023). Additionally, to mitigate the computational costs of training and deploying large LLMs, we use a large batch size m for sparse updates.
Non-symmetric structure to balance exploitation and exploration. The framework also features a non-symmetric structure as we divide the two agents into a main agent and an enhancer.
@@
1:  Input: offline dataset $\mathcal{D}_{\mathrm{off}}$ (can be empty); batch size $m>0$, and preference model $\mathbb{P}$.
@@
4:     Exploration with the enhancer: define the uncertainty quantifier with respect to $\pi_{t}^{1}$ as $\Gamma^{m}_{t}(\lambda,\pi_{t}^{1},\pi^{2})$ and compute the policy $\pi_{t}^{2}=\mathop{\mathrm{argmax}}_{\pi^{2}\in\Pi_{t}}\Gamma^{m}_{t}(\lambda,\pi_{t}^{1},\pi^{2})$ where
$\Pi_{t}=\{\pi^{\prime}\in\Pi:\underbrace{\eta\mathbb{E}_{x\sim d_{0}}D_{\mathrm{KL}}(\pi(\cdot|x),\pi^{1}(\cdot|x))}_{\text{How far does the enhancer move away.}}\leq\underbrace{\Gamma^{m}_{t}(\lambda,\pi_{t}^{1},\pi^{\prime})}_{\text{How much information we can get.}}\}.$
@@
Theorem 1 (Informal Theorem 2 in (Xiong et al.,, 2023)).
For any precision parameter $\epsilon>0$, with a batch size $m=\widetilde{O}(\frac{d_{e}}{\epsilon^{2}})$ and other suitable choices of hyper-parameters, then, with high probability, after at most $T=\widetilde{O}(d_{e})$ iterations, we can find a $t_{0}\in[T]$ so that
$J(\pi^{*})-J(\pi_{t_{0}})+\eta\mathbb{E}_{x_{t_{0}}\sim d_{0}}\big{[}D_{\mathrm{KL}}(\pi^{*}(\cdot|x_{t_{0}})\|\pi_{t_{0}}(\cdot|x_{t_{0}}))\big{]}\lesssim\epsilon,$
where $J(\pi)=\mathbb{E}_{x\sim d_{0}}[\mathbb{E}_{a\sim\pi(\cdot|x)}[r^{*}(x,a)]-\eta D_{\mathrm{KL}}(\pi(\cdot|x)\|\pi_{0}(\cdot|x))]$ is the KL-regularized value. Here $\widetilde{O}$ hides some log factors and $d_{e}$ is a complexity measure of the RLHF problem. In particular, if the reward function can be embedded into a $d$-dimensional space that is linear in the feature map of $\phi:\mathcal{X}\times\mathcal{A}\to\mathbb{R}^{d}$, we have $d_{e}=d$. (Zhong et al.,, 2022; Liu et al., 2023b, )
@@ 3.3 Practical Implementation Details
The MLE policy. Since the main agent only exploits the data, we can run DPO on the historical data to approximate the optimal policy under the $r_{\mathrm{MLE}}$: $\pi_{t}^{\mathrm{MLE}}$. We remark that while we use DPO here due to its simplicity, the Algorithm 1 can be implemented by combining it with any oracle algorithms (e.g., PPO and InfoNCA (Chen et al., 2024a, )) that are a reasonable approximation of the KL-regularized optimization problem.
Exploration policy. The primary challenge lies in the choice of enhancer policy for exploration. Recall that our goal is to find an enhancer policy that maximizes the relative uncertainty to the main agent from the following policy subset:
$\widetilde{\pi}\in\Pi_{t}=\{\pi^{\prime}\in\Pi:\underbrace{\eta\cdot\mathbb{E}_{x\sim d_{0}}D_{\mathrm{KL}}(\pi(\cdot|x),\pi_{t}^{1}(\cdot|x))}_{\text{How far does the enhancer move away.}}\leq\underbrace{\Gamma^{m}_{t}(\lambda,\pi_{t}^{1},\pi^{\prime})}_{\text{How much information we can get.}}\}.$
@@
Unfortunately, the uncertainty estimator does not have an analytical form except the linear case. But the main insight we can derive here is to maximize the policy difference with $\pi_{t}^{1}$, while maintaining a moderate KL divergence. This motivates us to use model variants of $\pi_{t}^{1}$. We discuss some popular heuristic implementations here.
@@
  • Adjusting Temperature and Training Steps. In the project of Claude (Bai et al., 2022a, ), the authors choose to use the models with different training steps as $(\pi_{t}^{1},\pi_{t}^{2})$. For instance, if we run PPO for $2$ epoch in total, we may take $\pi_{t}^{1}$ as the model saved at the end of the first epoch and take $\pi_{t}^{2}$ as the one saved at the end of the second epoch. Additionally, the LLaMA-2 project (Touvron et al.,, 2023) adjusts the sampling temperature of $\pi_{t}^{1}$ to induce $\pi_{t}^{2}$. These modifications introduce diversity in the models and facilitate exploration.
@@
  • Rejection Sampling is another popular ensemble-based exploration approach (Nakano et al.,, 2021; Dong et al.,, 2023; Gulcehre et al.,, 2023). In the context of LLMs, it is typically restricted to the best-of-$n$ sampling. Specifically, we sample $n$ independent responses by $\pi_{t}^{1}$ for each prompt, and then use a preference/reward function to rank the responses and take the one with the highest reward as the final output. In other words, we take $\pi_{t}^{2}$ as the best-of-n variant of $\pi_{t}^{1}$. In this way, the $\pi_{t}^{2}$ enlarges the margins between $\pi_{t}^{1}$ and provides exploration. Meanwhile, in this case, the KL divergence between the two policies is upper bounded by $\log n-\frac{n-1}{n}$ and is usually far better than this conservative estimation (Beirami et al.,, 2024).
@@
In our experiments, we use the DPO to approximate the computational oracle and implement DPO with the open-source package TRL [^5] https://github.com/huggingface/trl. We run DPO with the reference model $\pi_{0}$ (the SFT model) on the historical data for 2 epochs to get the MLE policy $\pi^{\mathrm{MLE}}_{t}$. We use a cosine learning rate scheduler with a peak learning rate of 5e-7 and 0.03 warm-up ratio. We use a global batch size of 128 and use a KL coefficient of $\eta=0.1$. To accelerate training, we do not restart from $\pi_{0}$ at each iteration as in Bai et al., 2022a ; Xiong et al., (2023) but use the last-iteration model as the initial checkpoint and use $\pi_{0}$ as the reference model. In this way, the data used for training is the same as that of Bai et al., 2022a and Xiong et al., (2023) but is of a different order. We do not see performance regression with this choice, and it saves us for half of the training time.

### v3

@@ 2 Reward Modeling as Human Feedback Approximation
We present the details of preference model construction in this section, where we study both the reward modeling as MLE of the BT model and the general preference model. We consider two versions of the training set: mix1: HH-RLHF + SHP + UltraFeedback + Summarization, which is the mixture of dataset used to train the current state-of-the-art open-source model. (Stiennon et al., 2020), and mix2: all the open-source datasets we collect, with details provided in Table 5.
@@ 2.1 Bradley-Terry Reward Model and Preference Model
Bradley-Terry model construction. We follow the previous works (Ouyang et al., 2022; Bai et al., 2022a) to initialize the reward model using the SFT model [^2] For preference/reward modeling, we use meta-llama/Meta-Llama-3-8B-Instruct since only this checkpoint is available in the early stage of this project. . We replace the last layer with a linear head to predict a scalar score suitable for preference learning. The reward model is trained using the negative log-likelihood loss function, enabling maximum likelihood estimation (MLE). This loss function is defined as:
@@
where $a^{w}$ is the preferred response over $a^{l}$. We train the LLaMA-3-8B-based reward model for one epoch with a global batch size of 512. The learning rate is set to $\mathrm{lr}=2\times 10^{-6}$, and a cosine learning rate schedule with a warm-up ratio of 0.03 is employed.
Preference model construction. A (pairwise) preference model takes a prompt $x$ and two responses $a^{1},a^{2}$ as the input and predicts the probability of ^$\hat{\mathbb{P}}(a^{1}\succ a^{2}|x,a^{1},a^{2})$ (Jiang et al., 2023). We follow Zhao et al. (2023); Liu et al. (2023a) to leverage the LLM’s capability as a next-token predictor for preference modeling. Specifically, for a given preference pair $(x,a^{1},a^{2},A)$, where $A$ indicates that the first response is preferred, the pair is formatted as an instruction-following task:
@@
If the second response is preferred, we replace the label A with B. Then, we simply treat the preference modeling as an instruction-following task to fine-tune the model on these instruction-label pairs. To mitigate position bias (the preference model may prefer the response that is given in the position of RESPONSE A), the order of the responses is randomized during data formatting. During inference, if we denote the probability of decoding A as $p_{A}$ and the probability of decoding B as $p_{B}$, then ^$\hat{\mathbb{P}}(a^{1}\succ a^{2}|x,a^{1},a^{2})$ is taken as $p_{A}/(p_{A}+p_{B})$. We train the LLaMA-3-8B-based preference model for one epoch. The samples are packed into blocks with length 3072 and a global batch size of 128 is used. The learning rate is set to $\mathrm{lr}=5\times 10^{-6}$, and a cosine learning rate schedule with a warm-up ratio of 0.03 is employed. We mention in passing that it is possible to include detailed rubrics in the data format to further improve the preference dataset, which we leave for future work (Qin et al., 2023).
@@ 2.2 Evaluation Result
We evaluate the models using the RewardBench (Lambert et al., 2024), a benchmark designed to assess reward model capabilities in four categories: Chat, Chat-Hard, Safety, and Reasoning. The main evaluation results are in Table 1. It is evident that without explicit training, the prompting approach is inferior to both the BT model and the preference model across all metrics. Meanwhile, the preference model outperforms the BT model in reasoning tasks related to coding and math. We also notice that with more data, especially data specified in coding, math, and safety, the reward model trained by mix2 achieves higher accuracy in safety and reasoning compared with early versions of attempts. In particular, we use the Ultra-RM-13B as a reference model in Table 1, where we can observe that the extra data related to safety and reasoning (as well as a stronger base model) largely contribute to the superior performance of our reward model.
Length bias in reward modeling. It is known that the LLMs aligned by RLHF usually give longer responses (Xiong et al., 2023; Yuan et al., 2024b; b), where the length bias also exists in the reward models, likely influenced by the preference data used. To better understand this bias, we randomly sample 2K prompts from the prompt set and use the SFT model to generate 8 responses per prompt. Then, we compute the lengths and rewards of the responses and plot the heatmaps of the Pearson correlation coefficient between them in Figure 3. Clearly, both of the two reward models are biased toward the longer responses to some degree. In comparison, UltraRM-13B demonstrates a stronger bias, as we observe that the mean coefficient of the UltraRM-13B (left) is 0.19, while it is 0.06 for our BT reward (right). This may partially result from the use of additional Capybara, OpenOrca, and UltraInteract, whose preferred responses are shorter than the rejected responses. We will return to the ablation study of the impacts of the reward models in Section 4.
Figure 3: The heatmap of the Pearson correlation coefficients between reward and response length. For each prompt, we use the SFT model to generate 16 responses and compute the coefficient. We also group the prompts by the mean responses (y-axis). The left figure is the UltraRM-13B and the right one is our BT reward based on mix2. We observe that the heatmaps for both UltraRM-13B and our BT reward model show a tendency toward the positive side. This indicates that the reward models exhibit a preference for longer responses.
Table 1: Comparison of the test accuracy between the Bradley-Terry (BT) reward model and the preference model. We evaluate the model using the Reward-Bench (Lambert et al., 2024).
@@ 3 Iterative Policy Optimization
We develop the main algorithms for the online iterative RLHF in this section, with both theoretical insights and implementation details. In particular, the algorithms are in a direct preference learning style for stable and efficient training.
@@ 3.1 Supervised Fine-Tuning
The base model used in this project is LLaMA-3-8B. To ensure the reproducibility and openness of the project, we perform SFT by ourselves to obtain the initial policy $\pi_{0}$. We collect a set of high-quality instruction datasets for SFT, such as ShareGPT (Chiang et al., 2023), SlimOrca (Lian et al., 2023b), MathInstruct (Yue et al., 2023), and Evol-Instruct (Xu et al., 2023a) (see the Appendix for a full list). The training is carried out for one epoch with a learning rate of $2\times 10^{-5}$. A cosine scheduler is employed, and the global batch size is set to 32 with a warm-up ratio of 0.03. To accelerate training, we follow Diao et al. (2023); Tunstall et al. (2023) to pack the samples and use a block size of 8192.
@@ 3.2 Iterative Direct Preference Learning: Theoretical Insights and Algorithmic Principles
From a theoretical perspective, one of the most critical metrics in the reinforcement learning process is the uncertainty estimator $\Gamma$. For example, considering the linear case, $r=\langle\theta,\phi(x,a)\rangle$ such that $\theta\in\mathbb{R}^{d},\|\theta\|\leq B$. For any two policies $\pi_{t}^{1},\pi_{t}^{2}$, we define the information gain as $\Gamma_{t}(\lambda,\pi_{t}^{1},\pi_{t}^{2})=\beta\|\mathbb{E}_{\pi_{t}^{1}}\phi(x,a^{1}_{t})-\mathbb{E}_{\pi_{t}^{2}}\phi(x,a^{2}_{t})\|_{\Sigma_{t}^{-1}}$ ($\beta$ is a constant coefficient), which is the projection of the new feature difference to the historical feature covariance matrix $\Sigma_{t}=\lambda I+\sum_{s=1}^{t-1}\mathbb{E}_{x\sim d_{0},a^{1}\sim\pi_{s}^{1},a^{2}\sim\pi_{s}^{2}}\left[(\phi(x,a^{1})-\phi(x,a^{2}))(\phi(x,a^{1})-\phi(x,a^{2}))^{\top}\right]$.
@@
Hybrid batch learning. We formulate a slightly more general framework to combine an initial offline dataset with online data collected during training, hence the name hybrid learning, similar to the recipe of Claude (Bai et al., 2022a) and LLaMA-2 (Touvron et al., 2023). We also use a large batch size m for sparse updates.
Non-symmetric structure to balance exploitation and exploration. The framework also features a non-symmetric structure by involving a main agent and an enhancer.
@@
1:  Input: offline dataset $\mathcal{D}_{\mathrm{off}}$ (can be empty); batch size $m>0$, and preference model $\mathbb{P}$, the number of iteration T.
@@
4:     Exploration with the enhancer: given the policy, assume we have the uncertainty quantifier with respect to $\pi_{t}^{1}$ as $\Gamma^{m}_{t}(\lambda,\pi_{t}^{1},\pi^{2})$ and compute the policy $\pi_{t}^{2}=\mathop{\mathrm{argmax}}_{\pi^{2}\in\Pi_{t}}\Gamma^{m}_{t}(\lambda,\pi_{t}^{1},\pi^{2})$ where
$\Pi_{t}=\{\pi^{\prime}\in\Pi:\underbrace{\eta\mathbb{E}_{x\sim d_{0}}D_{\mathrm{KL}}(\pi(\cdot|x),\pi^{1}(\cdot|x))}_{\text{How far does the enhancer move away.}}\leq\underbrace{\Gamma^{m}_{t}(\lambda,\pi_{t}^{1},\pi^{\prime})}_{\text{How much information we can get.}}\}.$ (8)
@@
Theorem 1 (Informal Theorem 2 in (Xiong et al., 2023)).
For any precision parameter $\epsilon>0$, with a batch size $m=\widetilde{O}(\frac{d_{e}}{\epsilon^{2}})$ and other suitable choices of hyper-parameters, then, with high probability, after at most $T=\widetilde{O}(d_{e})$ iterations, we can find a $t_{0}\in[T]$ so that $J(\pi^{*})-J(\pi_{t_{0}})+\eta\mathbb{E}_{x_{t_{0}}\sim d_{0}}\big{[}D_{\mathrm{KL}}(\pi^{*}(\cdot|x_{t_{0}})\|\pi_{t_{0}}(\cdot|x_{t_{0}}))\big{]}\lesssim\epsilon,$ where $J(\pi)=\mathbb{E}_{x\sim d_{0}}[\mathbb{E}_{a\sim\pi(\cdot|x)}[r^{*}(x,a)]-\eta D_{\mathrm{KL}}(\pi(\cdot|x)\|\pi_{0}(\cdot|x))]$ is the KL-regularized value. Here $\widetilde{O}$ hides some log factors and $d_{e}$ is a complexity measure of the RLHF problem. In particular, if the reward function can be embedded into a $d$-dimensional space that is linear in the feature map of $\phi:\mathcal{X}\times\mathcal{A}\to\mathbb{R}^{d}$, we have $d_{e}=d$. (Zhong et al., 2022; Liu et al., 2023b)
@@ 3.3 Practical Implementation Details
The MLE policy. Since the main agent only exploits the data, we can run DPO on the historical data to approximate the optimal policy under the $r_{\mathrm{MLE}}$: $\pi_{t}^{\mathrm{MLE}}$. We remark that while we use DPO here due to its simplicity, the Algorithm 1 can be implemented by combining it with any oracle algorithms (e.g., PPO and InfoNCA (Chen et al., 2024a)) that are approximations of the KL-regularized optimization problem.
@@
Exploration policy. The primary challenge lies in the choice of enhancer policy for exploration. Recall that our goal is to find an enhancer policy that maximizes the relative uncertainty to the main agent from the confidence set defined in Equation (8). Unfortunately, the uncertainty estimator does not have an analytical form except the linear case. But the main insight we can derive here is to maximize the policy difference with $\pi_{t}^{1}$, while maintaining a moderate KL divergence. This motivates us to use model variants of $\pi_{t}^{1}$. We discuss some popular heuristic implementations here.
@@
  • Adjusting Temperature and Training Steps. In the project of Claude (Bai et al., 2022a), the authors choose to use the models with different training steps as $(\pi_{t}^{1},\pi_{t}^{2})$. For instance, if we run PPO for $2$ epoch in total, we may take $\pi_{t}^{1}$ as the model saved at the end of the first epoch and take $\pi_{t}^{2}$ as the one saved at the end of the second epoch. Additionally, the LLaMA-2 project (Touvron et al., 2023) adjusts the sampling temperature of $\pi_{t}^{1}$ to induce $\pi_{t}^{2}$. These modifications introduce diversity in the models and facilitate exploration.
@@
  • Rejection Sampling is another popular ensemble-based exploration approach (Nakano et al., 2021; Dong et al., 2023; Gulcehre et al., 2023). In the context of LLMs, it is typically restricted to the best-of-$n$ sampling. Specifically, we sample $n$ independent responses by $\pi_{t}^{1}$ for each prompt, and then use a preference/reward function to rank the responses and take the one with the highest reward as the final output. In other words, we take $\pi_{t}^{2}$ as the best-of-n variant of $\pi_{t}^{1}$. In this way, the $\pi_{t}^{2}$ enlarges the margins between $\pi_{t}^{1}$ and provides exploration. Meanwhile, in this case, the KL divergence between the two policies is upper bounded by $\log n-\frac{n-1}{n}$ and is usually far better than this conservative estimation (Beirami et al., 2024);
  • After the first submission of this work, there is a line of works proposing to use biased DPO loss in the online iterative framework to encourage exploration (Xie et al., 2024; Zhang et al., 2024; Cen et al., 2024), whose idea originates from the theoretical RL study (Xiong, 2023; Liu et al., 2023b). Specifically, they add a SFT loss term, also known as the “feel-good” term, into the loss function so that the algorithm favors the model that is more optimistic. We refer the interested readers to these works for a more detailed algorithm description and empirical results. We also integrate the option of adding such a bias term in the revision of our public code.
@@
In our experiments, we use the DPO to approximate the computational oracle and implement DPO with the open-source package TRL [^3] https://github.com/huggingface/trl. We run DPO with the reference model $\pi_{0}$ (the SFT model) on the historical data for 2 epochs to get the MLE policy $\pi^{\mathrm{MLE}}_{t}$. We use a cosine learning rate scheduler with a peak learning rate of 5e-7 and 0.03 warm-up ratio. We use a global batch size of 128 and use a KL coefficient of $\eta=0.1$. To accelerate training, we do not restart from $\pi_{0}$ at each iteration as in Bai et al. (2022a); Xiong et al. (2023) but use the last-iteration model as the initial checkpoint and use $\pi_{0}$ as the reference model. In this way, the data used for training is the same as that of Bai et al. (2022a) and Xiong et al. (2023) but is of a different order. We do not see performance regression with this choice, and it saves us for half of the training time.
Algorithm 2 Practical Version of Online Iterative RLHF with BT Reward Model
1:  Input: offline dataset $\mathcal{D}_{\mathrm{off}}$ (can be empty); batch size $m>0$, rejection sampling parameter $n$ and reward model $r$; the number of iteration T.
2:  for t=1,…,T do
3:     Compute $\pi_{t}$ by the DPO algorithm with $\mathcal{D}_{\mathrm{off}}\cup\mathcal{D}_{1:t-1}$ using the SFT policy $\pi_{0}$ as reference model.
4:     Sample a batch of prompts $\{x_{i}\}_{i=1}^{m}$ from $d_{0}$. For each prompt, we sample $n/2$ responses using $\pi_{t}$ with temperature $1.0$ and $n/2$ responses using $\pi_{t}$ with temperature $0.7$.
5:     For each prompt $x_{i}$, we rank them using $r$ and take the best response and the worst one to construct a preference pair into $\mathcal{D}_{t}$. Eventually, we collect $m$ preference pairs.
6:  end for
7:  Output: the best policy in $(\pi^{1}_{1:T})$ by a validation set.